MATH 351 – Modern Algebra Problem Set Guidelines Fall 2018

Weekly homework assignments will be posted each Friday on the course homepage and collected the following Friday. These assignments will usually consist 10-12 problems (depending on level of difficulty, how many parts there are to each question, and so forth). Here are some guidelines to follow while doing your assignments, for our mutual benefit:

• Please be neat! Write up the problems in order. Use only on one side of the page (don’t worry about saving paper), and staple your assignment together.

• Leave lots of space for me to write comments.

• Copy out the problem statement before giving the solution. (This should help you make sure you have answered everything in the question.)

• Show enough detail so that a confused student in the class could follow your solution. Don’t expect that the reader already knows how to solve the problem. Just writing an answer is almost never sufficient.

• Proofs should be written in complete sentences, although your sentences will often contain mathematical symbols and statements. Proofread what you have said to be sure it makes sense.

• No “and then a miracle occurs” moments during proofs, please! If you are left with a gap in the proof that you just can’t bridge, acknowledge it. I will indicate how to fill the gap, if appropriate, in my comments.

• Start the assignment early (i.e. the day it is posted). This will allow you ample time to consult with me or your classmates if you get stuck on some problems. If you start the assignment on the day before it is due, you will be very unhappy!

• Work on making your arguments clear and concise. Make good use of notation and diagrams. As the semester progresses, I will grade your proofs more and more for both accuracy and presentation.

• If you work with other students to solve a problem, you must write up your final solutions independently. (Plan on taking scrap notes when you work out the problem first, and then “writing it up” in a clean form by yourself afterwards.) Add a note to your solution listing the other people you consulted.